### 19 Apr 2026 v3

# Thomassen’s proof and Filippov’s proof of the Weak Jordan Theorem

Abstract
We present, in detail and with a modern rigor, the two title proofs. The Weak Jordan Theorem (WJT) states that the complement of any topological circuit in the plane is disconnected.
1 JT, WJT and AT. Auxiliary results
Let us begin with a few necessary definitions. We denote by $I$ any compact real interval $I=[a,b]$ with $a<b$. If $f\colon I\to X$ is a map, we set $f[I]=\{f(x)\colon\;x\in I\}$, $f^{0}=\{f(x)\colon\;a<x<b\}$ and $\mathrm{ep}(f)=\{f(a),f(b)\}$. We also say that $f$ joins $f(a)$ to $f(b)$ in $X$. If $f(a)=f(b)$, we call $f$ closed. If $g\colon A\to B$ is any map and $C$ is any set, we define
$g[C]=\{g(x)\colon\;x\in A\cap C\}\,\text{ and }g^{-1}[C]=\{x\in A\colon\;g(x)\in C\}\,.$
An arc is a continuous injection $f\colon I\to\mathbb{R}^{2}$. A circuit is a continuous map $f\colon I=[a,b]\to\mathbb{R}^{2}$ such that $f(x)=f(y)$ for $x\neq y$ if and only if $\{x,y\}=\{a,b\}$. The following Jordan Theorem, or JT, is a fundamental result in plane topology. It is due to C. Jordan [14, pp. 587–594] in 1887.
Theorem 1.1 (Jordan, 1887)
For any circuit $f\colon I\to\mathbb{R}^{2}$, the complement
$\mathbb{R}^{2}\setminus f[I]=A\cup B$
is a disjoint union of two nonempty open connected sets $A$ and $B$ such that $A$ is bounded, $B$ is unbounded, and $\partial A=\partial B=f[I]$.
We explain the statement. We work in the Euclidean plane, which is a metric space $\langle\mathbb{R}^{2},d_{2}\rangle$ with the metric
$d_{2}(a,\,b)=\sqrt{(a_{x}-b_{x})^{2}+(a_{y}-b_{y})^{2}}\,$,
and in its subspaces. We count among them $\mathbb{R}$ and $I$. A set $X\subset M$ in a metric space $\langle M,d\rangle$ is disconnected if there exist open (equivalently, closed) sets $A,B\subset M$ such that
$A\cup B\supset X\;\&\;A\cap X\neq\emptyset\neq B\cap X\;\&\;A\cap B\cap X=\emptyset$
— we say that $A$ and $B$ cut $X$. Else, if such sets $A$ and $B$ do not exist, $X$ is connected. We take for granted that every real interval (not necessarily compact) is connected. A set $X\subset M$ is bounded if $X\subset B=B(a,r)$ for a ball $B$ in $\langle M,d\rangle$. Else, if $X$ is not contained in any ball, $X$ is unbounded. Finally, the boundary of $X$ is the set
$\partial X=\{a\in M\colon\;\forall r\,B(a,r)\cap X\neq\emptyset\neq B(a,r)\cap(M\setminus X)\}\,.$
In this article, we will not prove the JT, but we present two proofs for the weaker version, the Weak Jordan Theorem, or WJT.
Theorem 1.2 (WJT)
For any circuit $f\colon I\to\mathbb{R}^{2}$, the complement
$\mathbb{R}^{2}\setminus f[I]$
is disconnected.
JT $\Rightarrow$ WJT because the disjoint union of two nonempty open sets is disconnected. The proof of the JT is harder than that of the WJT because it requires the following “lemma”, the Arc Theorem, or AT.
Theorem 1.3 (AT)
For any arc $f\colon I\to\mathbb{R}^{2}$, the complement
$\mathbb{R}^{2}\setminus f[I]$
is connected.
We write about this overlooked cornerstone of the theory of planar graphs in a separate contribution [15], where we also complete the presentation of a proof of the JT.
In 1992, C. Thomassen found in [17] an interesting proof of the JT based on planar graphs. He actually obtained combinatorial proofs of the more general Jordan–Schönflies theorem and of the theorem on classification of surfaces. In Section 2 we present only a part of [17], Thomassen’s proof of the WJT. It employs the following virtual configurations. A $K_{3,3}$-configuration has two disjoint three-element sets $U,V\subset\mathbb{R}^{2}$ and nine arcs $f_{u,v}\colon I\to\mathbb{R}^{2}$ for $u\in U$ and $v\in V$ such that $f_{u,v}$ joins $u$ to $v$ and
$f_{u,\,v}^{0}\cap f_{u^{\prime},\,v^{\prime}}^{0}=\emptyset=f_{u,\,v}^{0}\cap(U\cup V)$
for every two pairs $\langle u,v\rangle\neq\langle u^{\prime},v^{\prime}\rangle$. It is a plane drawing, without crossings, of the complete bipartite graph $K_{3,3}$; these configurations do not exist. A PL map is a map $f\colon[a,b]\to\mathbb{R}^{2}$ such that for a partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of $[a,b]$, all restrictions $f\,|\,[a_{i-1},a_{i}]$ for $i=1,2,\dots,k$ are non-constant linear maps. Every PL map is continuous. The straight segments $f[\,[a_{i-1},a_{i}]\,]$ ($\subset\mathbb{R}^{2}$) are the segments of $f[I]$. The points $f(a_{i})$ ($\in\mathbb{R}^{2}$) are the corners of $f[I]$. In a PL $K_{3,3}$-configuration, the nine arcs are PL maps. Thomassen’s proof of the WJT in [17], presented in Section 2, is the sequence claim 1 $\Rightarrow$ claim 2 $\Rightarrow$ claim 3 $\Rightarrow$ claim 4, where the four claims are as follows.
  1. 1.
    The JT holds for PL circuits.
  2. 2.
    PL $K_{3,3}$-configurations do not exist.
  3. 3.
    $K_{3,3}$-configurations do not exist.
  4. 4.
    The WJT holds.
Thomassen’s main invention is the last implication claim 3 $\Rightarrow$ claim 4.
Another proof of the WJT is contained in the proof of the JT given by A. F. Filippov in [5] in 1950. We present this proof in Section 3. It can be divided in three steps.
  1. 1.
    Let $f,g\colon I=[a,b]\to\mathbb{R}^{2}$ be PL maps such that $f[I]\cap g[I]=\emptyset$ and that either $f$ is closed or $g[I]$ lies between the two vertical lines going through the points $f(a)$ and $f(b)$. Consider the function
    $N=N(c)=N(c,\,f)\colon g[I]\to\{0,1\}$
    such that $N(c)$ is the parity of the number of simple intersections of the vertical half-line going up from the point $c$ with $f[I]$ (we precisely define $N$ later). Then $N$ is constant.
  2. 2.
    The function $N(c,f)$ is additive in the variable $f$. If $f$ is not closed and the point $c$ lies between the two vertical lines and below all intersections of the line $x=c_{x}$ with $f[I]$, then $N(c)=1$.
  3. 3.
    Let $f\colon I\to\mathbb{R}^{2}$ be a circuit. The results in steps 1 and 2 are used to obtain two distinct points $c,d\in\mathbb{R}^{2}\setminus f[I]$ such that for every PL map $g$ joining $c$ to $d$ we have $g[I]\cap f[I]\neq\emptyset$. This means that the complement $\mathbb{R}^{2}\setminus f[I]$ is disconnected and the WJT holds.
Why this presentation of the two proofs? They are beautiful, but they are not rigorous by modern standards. Since we like them, we decided to elevate them to the present level of rigor. See also the discussion in Section 4.
Auxiliary results. In the rest of Section 1 we present, with proofs, auxiliary results that will be needed in Sections 2 and 3. Since they are mostly standard, the impatient reader may now jump to the two proofs and check any auxiliary result only if the need arises.
We begin with three constructions related to maps. Let $I=[a,b]$ and $J=[c,d]$ be real intervals with $a<b$ and $c<d$, and let $f\colon I\to X$ be a map. We can parametrize $f[I]$ by $J$ via the map $g=f(h)\colon J\to X$, where $h\colon J\to I$ is the linear homeomorphism $h(x)=\frac{b-a}{d-c}(x-c)+a$. Clearly, $g$ joins $g(c)=f(a)$ to $g(d)=f(b)$ in $X$. If $f$ is a PL map, then so is $g$, and the same holds for being a continuous map, an arc, or a circuit. Thus we can use any compact interval $I=[\alpha,\beta]$, $\alpha<\beta$, as the common definition domain of PL maps, (PL) arcs, and (PL) circuits.
Let $f,g\colon I=[a,b]\to X$ be two maps such that $f(b)=g(a)$. We write $f+g$ for their concatenation $h=f+g\colon I\to X$ given by h(x)=f(2xa)h(x)=f(2x-a) for $a\leq x\leq\frac{1}{2}(a+b)$, and by h(x)=g(2xb)h(x)=g(2x-b) for $\frac{1}{2}(a+b)\leq x\leq b$. Clearly, $h$ joins $h(a)=f(a)$ to $h(b)=g(b)$ in $X$. If $f$ and $g$ are PL maps, then so is $h$, and the same holds for continuity. Let in addition $f[I]\cap g[\,(a,b]\,]=\emptyset$, and let $f$ and $g$ be arcs. Then $f+g$ is an arc. Let in addition $f[I]\cap g^{0}=\emptyset$, $g(b)=f(a)$, and let $f$ and $g$ be arcs. Then $f+g$ is a circuit.
Finally, let $U$, $V$, and $\{f_{u,v}\colon\;u\in V,v\in V\}$ be a $K_{3,3}$-configuration. For any of the arcs $f_{u,v}\colon I=[a,b]\to\mathbb{R}^{2}$, we denote by $f_{v,u}\colon I\to\mathbb{R}^{2}$ the reverse arc
$f_{v,\,u}(x)=f_{u,\,v}(a+b-x)\,.$
The first two auxiliary results concern closed (and non-compact) sets in metric spaces used in Filippov’s proof, such as vertical half-lines in $\mathbb{R}^{2}$. Recall that a set $X\subset M$ in a metric space $\langle M,d\rangle$ is closed if the complement $M\setminus X$ is an open set, that is, if for every point $a\in M\setminus X$ there exists an $r>0$ such that the ball $B(a,r)\subset X\setminus M$.
Proposition 1.4
In any metric space, the union of two closed sets is a closed set.
Proof. Let $X,Y\subset M$ be closed sets in a metric space $\langle M,d\rangle$, and let
$a\in M\setminus(X\cup Y)=(M\setminus X)\cap(M\setminus Y)\,.$
We have $B(a,r)\subset M\setminus X$ and $B(a,r^{\prime})\subset M\setminus Y$ for some $r,r^{\prime}>0$. Thus
$B(a,\,r^{\prime\prime})\subset(M\setminus X)\cap(M\setminus Y)$
for $r^{\prime\prime}=\min(r,r^{\prime})$ (>0>0). We see that $X\cup Y$ is closed. $\Box$
Proposition 1.5
Let $\ell\subset\mathbb{R}^{2}$ be a vertical half-line going up or down from a point $c\in\mathbb{R}^{2}$. Then $\ell$ is a closed set.
Proof. Suppose that $\ell$ goes down (if it goes up, we argue similarly) and that $a\in\mathbb{R}^{2}\setminus\ell$. If $a_{x}=c_{x}$, we define $r=a_{y}-c_{y}$ (>0>0), and if $a_{x}\neq c_{x}$, we set $r=|a_{x}-c_{x}|$ (>0>0). Then $d_{2}(a,b)\geq r$ for every point $b\in\ell$, and therefore $B(a,r)\subset\mathbb{R}^{2}\setminus\ell$. $\Box$
The second group of auxiliary results concerns compactness in metric spaces. Recall that a set $X\subset M$ in a metric space $\langle M,d\rangle$ is compact if every sequence in $X$ has a convergent subsequence with a limit in $X$. We take for granted the compactness of $I=[a,b]$ ($\subset\mathbb{R}^{2}$).
Proposition 1.6
Let $\langle M,d\rangle$ and $\langle N,e\rangle$ be metric spaces, $f\colon M\to N$ be a continuous map, and let $X\subset M$ be a compact set. Then $f[X]$ is a compact subset of $N$.
Proof. Let $(b_{n})\subset f[X]$ be a sequence. Using the axiom of choice, we select elements $a_{n}\in X$ such that $f(a_{n})=b_{n}$. By the compactness of $X$, there is a subsequence $(a_{m_{n}})$ of $(a_{n})$ with $\lim a_{m_{n}}=a\in X$. Since $f$ is continuous,
$\lim b_{m_{n}}=\lim f(a_{m_{n}})=f(\lim a_{m_{n}})=f(a)\in f[X]\,.$
$\Box$
Proposition 1.7
Let $\langle M,d\rangle$ be a metric space and $X\subset M$ be a compact set. Then $X$ is a closed and bounded set.
Proof. We show that if $X$ is not closed or not bounded, then it is not compact. If $X$ is not closed, then there is a sequence $(a_{n})\subset X$ such that $\lim a_{n}=a\in M\setminus X$. Such $(a_{n})$ does not have any subsequence with a limit in $X$. If $X$ is not bounded, then there exists a sequence $(a_{n})\subset X$ such that $d(a_{m},a_{n})\geq 1$ for every $m<n$. Such $(a_{n})$ does not have any convergent subsequence. $\Box$
Proposition 1.8
For any continuous map $f\colon I\to\mathbb{R}^{2}$, the complement $\mathbb{R}^{2}\setminus f[I]$ is an open set.
Proof. By Proposition 1.6, the image $f[I]$ is a compact subset of $\mathbb{R}^{2}$. By Proposition 1.7, this image is a closed set. $\Box$
Proposition 1.9
Let $\langle M,d\rangle$ be a metric space, $X\subset M$ be a closed set, and let $f\colon I=[a,b]\to M$ be a continuous map. If $f[I]\cap X\neq\emptyset$, then the subset $f^{-1}[X]$ of $I$ has both minimum and maximum.
Proof. Since the nonempty set $f^{-1}[X]$ is closed and bounded, $\inf(f^{-1}[X])$ and $\sup(f^{-1}[X])$ are elements of it. $\Box$
Recall that two sets $X,Y\subset M$ in a metric space $\langle M,d\rangle$ have distance
$d(X,\,Y)=\inf(\{d(x,\,y)\colon\;x\in X,\,y\in Y\})\ \ (\in[0,\,+\infty))\,.$
For $X=\{x\}$ we write just $d(x,Y)$ instead of $d(\{x\},Y)$.
Proposition 1.10
Let $\langle M,d\rangle$ be a metric space and let $X,Y\subset M$ be nonempty compact sets that are disjoint. Then $d(X,Y)>0$.
We omit the proof because we prove a stronger result.
Proposition 1.11
Let $\langle M,d\rangle$ be a metric space and let $X,Y\subset M$ be nonempty sets such that $X$ is compact, $Y$ is closed and $X\cap Y=\emptyset$. Then $d(X,Y)>0$.
Proof. For the contrary, let $d(X,Y)=0$. Then there exist sequences $(x_{n})\subset X$ and $(y_{n})\subset Y$ such that $\lim d(x_{n},y_{n})=0$. Since $X$ is compact, we may assume that $\lim x_{n}=x\in X$. Using the triangle inequality, we see that also $\lim y_{n}=x$. Since $Y$ is closed, $x\in Y$. Thus $x\in X\cap Y$, which contradicts the disjointness of $X$ and $Y$. $\Box$
In Thomassen’s proof, Proposition 1.10 suffices. Filippov’s proof uses also Proposition 1.11.
Recall that a map $f\colon M\to N$ between metric spaces $\langle M,d\rangle$ and $\langle N,e\rangle$ is continuous at a point $a\in M$ if for every $\varepsilon>0$ there is $\delta>0$ such that $f[B(a,\delta)]\subset B(f(a),\varepsilon)$. The map $f$ is continuous if it is continuous at every point in $M$. The map $f$ is uniformly continuous if for every $\varepsilon>0$ there is $\delta>0$ such that $d(x,y)\leq\delta$ $\Rightarrow$ $e(f(x),f(y))\leq\varepsilon$ for every $x,y\in M$. It is easy to show that for every continuous map $f$ and every open (closed) set $X\subset N$ the preimage $f^{-1}[X]$ is an open (closed) set in $M$. If $f$ is continuous at $a$ then for every sequence $(a_{n})\subset M$ with $\lim a_{n}=a$ we have $\lim f(a_{n})=f(a)$.
Proposition 1.12
Let $\langle M,d\rangle$ and $\langle N,e\rangle$ be metric spaces, the former being compact, and let $f\colon M\to N$ be a continuous map. Then $f$ is uniformly continuous.
Proof. Suppose, for the contrary, that $f$ is not uniformly continuous. Then there exist $\varepsilon>0$ and sequences $(a_{n}),(b_{n})\subset M$ such that $\lim d(a_{n},b_{n})=0$, but $e(f(a_{n}),f(b_{n}))\geq\varepsilon$ for every $n$. Due to the compactness of $M$ we may assume that $\lim a_{n}=\lim b_{n}=a$ ($\in M$). But $\lim f(a_{n})=\lim f(b_{n})=f(a)$ ($\in N$) does not hold, in contradiction with the continuity of $f$ at $a$ $\Box$
The third group of auxiliary results concerns connectedness.
Proposition 1.13
Let $\langle M,d\rangle$ and $\langle N,e\rangle$ be metric spaces, $X\subset M$ be a connected set and $f\colon M\to N$ be a continuous map. Then $f[X]$ is a connected subset of $N$.
Proof. We show that if $f[X]$ is disconnected, then so is $X$. Let $A,B\subset N$ be open sets cutting $f[X]$. Since $f$ is continuous, the subsets $f^{-1}[A]$ and $f^{-1}[B]$ of $M$ are open. It is easy to see that they cut $X$. $\Box$
[…]
Proposition 1.14
A graph $G=\langle V,E\rangle$ is disconnected if and only if there exists a partition $\{A,B\}$ of $V$ such that no edge $e\in E$ joins $A$ and $B$.
Proof. Suppose that $G$ is disconnected and that $u,v\in V$ cannot be joined in $G$ by any path. We set
[…]
Then $u\in A$, $v\in B$ and no edge $e\in E$ joins $A$ and $B$. So $\{A,B\}$ is the desired partition of $V$.
Suppose that $A$ and $B$ are as stated, and take any $u\in A$ and $v\in B$. Any path in $G$ joining $u$ and $v$ would contain an edge $e\in E$ joining $A$ and $B$, which means that there is no such path. We see that $G$ is disconnected. $\Box$
The intersection graph corresponding to a set system $X_{v}$, $v\in V$, is the graph $\langle V,E\rangle$ with the edges $E=\{\{u,v\}\colon\;u,v\in V,u\neq v,X_{u}\cap X_{v}\neq\emptyset\}$.
Proposition 1.15
Let $\langle M,d\rangle$ be a metric space, let $X_{v}\subset M$, $v\in V$, be a set system of connected sets and let the intersection graph $G=\langle V,E\rangle$ be connected. Then the set
$X=\bigcup_{v\in V}X_{v}\ \ (\subset M)$
is connected.
Proof. Suppose that $X$ is disconnected and that the sets $A,B\subset M$ cut $X$. Every set $X_{v}$ is completely contained either in $A$ or in $B$, and we get the partition $\{V_{A},V_{B}\}$ of $V$ by the sets
$V_{A}=\{v\in V\colon\;X_{v}\subset A\}\,\text{ and }\,V_{B}=\{v\in V\colon\;X_{v}\subset B\}\,.$
Since $A\cap B\cap X=\emptyset$, there is no edge $e\in E$ joining $V_{A}$ and $V_{B}$. Hence $G$ is disconnected by Proposition 1.14. $\Box$
We prove, as an application, that balls in the plane are connected.
Proposition 1.16
Every ball in the metric space $\langle\mathbb{R}^{2},d_{2}\rangle$ is connected.
Proof. Let $B=B(b,r)=\{a\in\mathbb{R}^{2}\colon\;d_{2}(a,b)<r\}$ be a ball in the Euclidean plane. For any point $a\in B\setminus\{b\}$ we denote by $S_{a}$ the straight segment joining $b$ and $a$. Every $S_{a}$ is connected by Proposition 1.13, as a continuous image of an interval. Thus
$B=\bigcup_{a\in B\setminus\{b\}}S_{a}$}
is connected by Proposition 1.15 (the intersection graph is complete). $\Box$
In general, balls in metric spaces need not be connected.
The following proposition is the most intricate of the auxiliary results. We need a lemma for it.
Lemma 1.17
Let $X\subset\mathbb{R}^{2}$ be an open set, let $u,v\in X$ with $u\neq v$, and let there be a PL arc joining $u$ to $v$ in $X$. Then there exists a $v$-centered ball $B\subset X$ such that for every point $w\in B$ there exists a PL arc joining $u$ to $w$ in $X$.
Proof. Let $f\colon I=[a,b]\to X$ be a PL arc joining $u$ to $v$ in $X$, let $a=a_{0}<a_{1}<\dots<a_{k}=b$ be the partition of $[a,b]$, and let $s_{i}=f[\,[a_{i-1},a_{i}]\,]$ for $i=1,2,\dots,k$ be the segments of $f[I]$. Since $d_{2}(s_{i},v)>0$ for $i=1,2,\dots,k-1$ (Propositions 1.6 and 1.10), we can take small enough $r>0$ such that $B=B(v,r)\subset X$ and $s_{i}\cap B=\emptyset$ for $i=1,2,\dots,k-1$. We define for every point $w\in B$ a PL arc $f_{w}$ joining $u$ to $w$ in $X$. If $w\in B\cap s_{k}$, then $f_{w}$ is the initial part of $f$ ending at $w$. If $w\in B\setminus s_{k}$, then $f_{w}$ is the extension of $f$ by the straight segment joining $v$ to $w$. It is easy to see that $f_{w}$ is as claimed. $\Box$
Proposition 1.18
Let $X\subset\mathbb{R}^{2}$ be an open connected set and let $u,v\in X$ with $u\neq v$. Then there exists a PL arc joining $u$ to $v$ in $X$.
Proof. Let $Y\subset X$ be the set of points $x\in X$ such that $x=u$ or a PL arc joins $u$ to $x$ in $X$. Clearly, $Y\neq\emptyset$ ($u\in Y$) and $Y$ is open (Lemma 1.17). We show that also the set $Z=X\setminus Y$ is open. If $Z\neq\emptyset$, we have the contradiction that $X$ is disconnected. Thus $Z=\emptyset$ and $Y=X$, which implies that there exists a PL arc joining $u$ to $v$ in $X$.
Let $y\in Z$. We take a ball $B=B(y,r)\subset X$ such that $u\not\in B$, which is possible as $y\neq u$, and show that $B\subset Z$. Let $z\in B$ with $z\neq y$. Clearly, $z\neq u$. Suppose that $f\colon I\to X$ is a PL arc joining $u$ to $z$ in $X$. Let $s=yz$ ($\subset B$) be the straight segment joining $y$ and $z$. Considering the first intersection in $f[I]\cap s$ (Proposition 1.9), we obtain a PL arc joining $u$ to $y$ in $X$, but this contradicts $y\in Z$. Thus $f$ does not exist and $z\in Z$. We see that $B\subset Z$. The set $Z$ is open. $\Box$
The last group of auxiliary results is related to Thomassen’s construction of a plane drawing of $K_{3,3}$ based on a circuit whose complement is connected.
Proposition 1.19
Let $X\subset\mathbb{R}^{2}$ be a compact set. Then there exist four points $u,v,r,s\in X$ such that $u_{x}$ and $v_{x}$ have, respectively, the maximum and minimum $x$-coordinate on $X$, and $r$ and $s$ have an analogous property in the $y$-coordinate.
Proof. We prove the existence of $v$ (the existence of the other three points is established similarly). Since $X$ is bounded (Proposition 1.7), there exists a sequence $(a_{n})\subset X$ such that
$\lim_{n\to\infty}(a_{n})_{x}=\inf(\{a_{x}\colon\;a\in X\})=:\alpha\,.$
Since $X$ is compact, we may assume that $\lim a_{n}=v\in X$. Since the function $c\mapsto c_{x}$ is continuous,
$v_{x}=\big(\lim_{n\to\infty}a_{n}\big)_{x}=\lim_{n\to\infty}(a_{n})_{x}=\alpha$
and $v$ is a leftmost point of $X$. $\Box$
Proposition 1.20
Let $f\colon I=[a,b]\to\mathbb{R}^{2}$ be a circuit. Then there exist four points $u,v,r,s\in f[I]$ such that $u_{x}<v_{x}$ and $r_{y}<s_{y}$.
Proof. We prove the result for $u$ and $v$; the argument for $r$ and $s$ is similar. If $u$ and $v$ did not exist, then $f[I]$ would be a vertical straight segment $\{e\}\times[c,d]$. Let $f(a)=f(b)=\langle e,\alpha\rangle$ and $f(\frac{1}{2}(a+b))=\langle e,\beta\rangle$. Then $\alpha\neq\beta$, for example $\alpha>\beta$. By the intermediate value theorem, for every $\gamma\in(\beta,\alpha)$ there exist numbers $x\in(a,\frac{1}{2}(a+b))$ and $y\in(\frac{1}{2}(a+b),b)$ such that
$f(x)=f(y)=\langle e,\,\gamma\rangle\,.$
This contradicts the almost injectivity of $f$. $\Box$
Proposition 1.21
Let $f\colon I\to\mathbb{R}^{2}$ be a circuit and $u,v\in f[I]$ with $u\neq v$. Then there exist arcs $g,h\colon I\to\mathbb{R}^{2}$ such that $g^{0}\cap h^{0}=\emptyset$, $g[I]\cup h[I]=f[I]$, and $\mathrm{ep}(g)=\mathrm{ep}(h)=\{u,v\}$.
Proof. Wlog $f(a)=u$ and $f(b)=v$, where $a<b$ are in $I=[c,d]$. We define g=f|[a,b]g=f\,|\,[a,b] and $h\colon[b,d-c+a]\to\mathbb{R}^{2}$ by
$\text{$h(x)=f(x)$ for $b\leq x\leq d$ and $h(x)=f(x+c-d)$ for $d\leq x\leq d-c+a$}\,.$
According to the initial remark, we may redefine $g$ and $h$ on $I$. It is easy to see that $g$ and $h$ are as stated. $\Box$
Proposition 1.22
Let $f\colon I=[a,b]\to\mathbb{R}^{2}$ be continuous, $f(a)_{y}<f(b)_{y}$, and let $\ell$ be a horizontal line with the $y$-coordinate in $(f(a)_{y},f(b)_{y})$. Then $\ell\cap f[I]\neq\emptyset$.
Proof. We apply the intermediate value theorem to the continuous function $f(t)_{y}\colon I\to\mathbb{R}$. $\Box$
Proposition 1.23
Let $A,B\subset[c,d]$ be nonempty closed sets that are disjoint. Then there exists a subinterval $x,y]\subset[c,d]$ such that $x\in A$ and $y\in B$, or $x\in B$ and $y\in A$, and $(x,y)\cap(A\cup B)=\emptyset$.
Proof. Wlog we have numbers $a<b$ such that $a\in A$ and $b\in B$. We define $x=\sup(\{\alpha\in[a,\,b]\colon\;\alpha\in A\})$. Then $x\in A$ and $x<b$. We define $y=\inf(\{\alpha\in[x,\,b]\colon\;\alpha\in B\})$. Then $y\in B$ and $y>x$. Clearly, $(x,y)\cap(A\cup B)=\emptyset$. Thus $[x,y]$ is the desired subinterval. $\Box$
2 Thomassen’s proof of the WJT
Claim 1. The JT holds for PL circuits. Suppose that $f\colon I\to\mathbb{R}^{2}$ is a PL circuit. We first prove that the open set $D=\mathbb{R}^{2}\setminus f[I]$ (Proposition 1.8) is disconnected: we define a continuous map $g\colon D\to\mathbb{R}$ such that $g[D]=\{0,1\}$. By Proposition 1.13, the set $D$ is disconnected because $\{0,1\}$ is a disconnected subset of $\mathbb{R}$.
In order to define $g$ (we slightly simplify [17]), we take any point $a\in D$, draw a vertical half-line $\ell(a)$ up from $a$, and consider the geometric intersection $P(a)=\ell(a)\cap f[I]$. If $P(a)=\emptyset$, we set $g(a)=0$. Else
$P(a)=(z_{1}<z_{2}<\dots<z_{k}),\ k\geq 1\,$,
where << is the vertical order and every $z_{i}$ is either a union of consecutive (vertical) segments of $f[I]$ or a point in $f[I]$. The $z_{i}$ are mutually disjoint. We call $z_{i}$simple intersection if locally near $z_{i}$, the set $f[I]$ lies on both sides of $\ell(a)$. Else, if locally $f[I]$ lies on only one side of $\ell(a)$, we call $z_{i}$double intersection. We define $g(a)\in\{0,1\}$ as
[…]
We show that $g$ is continuous. Let $a\in D$. We take a ball $B=B(a,r)\subset D$ with radius $r>0$ so small that no half-line $\ell(a^{\prime})$ with $a^{\prime}\in B$ contains a corner of $f[I]$ not present in $P(a)$. If we move $a$ to $a^{\prime}\in B$, simple intersections in $P(a)$ are preserved, and every double intersection either does not change (when $\ell(a^{\prime})\subset\ell(a)$ or $\ell(a)\subset\ell(a^{\prime})$) or disappears or decays into two simple (point) intersections. Importantly, no new intersections besides those born from $P(a)$ are introduced. It follows that the number of simple intersections increases by an even number in $\mathbb{N}_{0}=\{0,1,\dots\}$, and therefore $g(a)=g(a^{\prime})$. The function $g$ is continuous at $a$ and on $D$.
Since $f[I]$ is a union of finitely many straight segments, it follows that $f[I]$ has a highest point, which is a corner. Considering the values of $g$ near this highest corner, we see that $g[D]=\{0,1\}$. This concludes the proof that $D$ is disconnected.
We prove that
$D=\mathbb{R}^{2}\setminus f[I]=L(f)\cup R(f)$
for two nonempty open connected and disjoint sets $L(f)$ and $R(f)$. To define them, we consider the partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of the domain $I=[a,b]$ of $f$ and list the corners $c_{i}=f(a_{i})$ and segments $s_{i}=f[\,[a_{i-1},a_{i}]\,]$ of $f[I]$ for $i=1,2,\dots,k$ in cyclic (indices are taken modulo $k$) order
$c_{0},\,s_{1},\,c_{1},\,s_{2},\,c_{2},\,s_{3},\,\dots,\,c_{k-1},\,s_{k},\,c_{k}=c_{0}\,.$
We fix one of the two orientations of $f[I]$ and define $L(f)$ as the set of points $c\in D$ that can be joined to $f[I]$ in $D$ from the left. It means that there is a PL map $f_{c}\colon I=[a,b]\to\mathbb{R}^{2}$ such that $f_{c}(a)=c$, $f_{c}(b)\in f[I]$, $f_{c}^{0}\subset D$, and the last segment of $f_{c}[I]$ approaches $f[I]$ from the left side with respect to the fixed orientation. Note that $f_{c}^{0}\subset L(f)$. We similarly define $R(f)$, by joining $c\in D$ to $f[I]$ in $D$ from the right. It is easy to see that the sets $L(f)$ and $R(f)$ are nonempty and open (Lemma 1.17 and its proof), and that $L(f)\cup R(f)=D$. If they are connected, it follows that they are disjoint: if $L(f)\cap R(f)\neq\emptyset$, then $L(f)\cup R(f)=D$ is connected by Proposition 1.15, which contradicts the previous result.
So it remains to prove that $L(f)$ and $R(f)$ are connected sets. The remaining properties stated in the JT, namely that one of the sets is bounded, the other is unbounded, and that $f[I]$ is the common boundary, follow easily.
We prove that $L(f)$ is connected (the argument for $R(f)$ is similar). We fix a point $c\in L(f)$ and for every point $d\in L(f)$ we define a connected set $C_{d}$ such that $c,d\in C_{d}\subset L(f)$. Then
$L(f)=\bigcup_{d\in D}C_{d}$}
is connected by Proposition 1.15 (the intersection graph is complete).
We define sets $C_{d}$. For $i=1,2,\dots,k$, we denote by $\ell_{i}$ ($\ni c_{i}$) the axis of the two angles determined by $s_{i}$ and $s_{i+1}$. For $\delta>0$ let $c_{i}(\delta)\in\ell_{i}$ be the point on $\ell_{i}$ lying in a distance of $\delta$ from $c_{i}$ on the left of $f[I]$. Let $s_{i}(\delta)$ be the straight segment joining $c_{i-1}(\delta)$ and $c_{i}(\delta)$. If $\delta$ is sufficiently small, then
$s_{i}(\delta)\cap s_{j}=\emptyset\,\text{ for }\,j=i-1,\,i,\,i+1\,.$
Also, $d_{2}(c,s_{i})\leq\delta$ for every $c\in s_{i}(\delta)$. It follows (Propositions 1.6 and 1.10) that there is a $\delta_{0}>0$ such that $s_{i}(\delta)\subset L(f)$ for every $\delta\leq\delta_{0}$. Let
$C(\delta)=\bigcup_{i=1}^{k}s_{i}(\delta)\,.$}
Then $C(\delta)\subset L(f)$ for every sufficiently small $\delta$. Every set $C(\delta)$ is connected by Proposition 1.13 (which implies that every segment $s_{i}(\delta)$ is connected) and Proposition 1.15 (the intersection graph spans a $k$-cycle). Let $c,d\in L(f)$ be the above mentioned points, and let $f_{c}$ and $f_{d}$ be respective PL maps that join them to $f[I]$ in $D$ from the left. We take small enough $\delta$ such that $C(\delta)\subset L(f)$ and that
$f_{c}^{0}\cap C(\delta)\neq\emptyset\neq f_{d}^{0}\cap C(\delta)\,.$
It follows that
$C_{d}=\{c\}\cup f_{c}^{0}\cup C(\delta)\cup f_{d}^{0}\cup\{d\}\ \ (\subset L(f))$
is the desired set. It is connected by Propositions 1.13 and 1.15.
Claim 2. PL $K_{3,3}$-configurations do not exist. We assume that $U$, $V$ ($\subset\mathbb{R}^{2}$) and $\{f_{u,v}\colon\;u\in U,v\in V\}$ is a PL $K_{3,3}$-configuration and obtain a contradiction. Let $U=\{u_{1},u_{3},u_{5}\}$ and $V=\{u_{2},u_{4},u_{6}\}$. We write $f_{i,j}$ instead of $f_{u_{i},u_{j}}$. We consider the PL circuit $C\colon I\to\mathbb{R}^{2}$ given by the $6$-cycle
$C=((((f_{1,\,2}+f_{2,\,3})+f_{3,\,4})+f_{4,\,5})+f_{5,\,6})+f_{6,\,1}\,.$
We fix one of the two orientations of $C[I]$ and consider the partition $D=\mathbb{R}^{2}\setminus C[I]=L(C)\cup R(C)$ of $D$ defined in Step 1. Let
$e_{1}=f_{1,\,4},\ e_{2}=f_{2,\,5}\,\text{ and }\,e_{3}=f_{3,\,6}$
be the three remaining arcs of the $K_{3,3}$-configuration. Each set $e_{i}^{0}$ lies completely either in $L(C)$ or in $R(C)$ ($e_{i}^{0}$ is connected by Proposition 1.13). Two of them, for example, $e_{2}^{0}$ and $e_{3}^{0}$, lie in one set, for example, $R(C)$. We consider the PL circuit $C^{\prime}\colon I\to\mathbb{R}^{2}$ given by the $4$-cycle
$C^{\prime}=((f_{2,\,3}+f_{3,\,4})+f_{4,\,5})+f_{5,\,2}\,.$
We orient $C^{\prime}[I]$ in accordance with the chosen orientation of $C[I]$ (the first three arcs of $C^{\prime}$ already have this orientation). Let $g=e_{3}+f_{6,5}$. We consider the partition
$D^{\prime}=\mathbb{R}^{2}\setminus C^{\prime}[I]=L(C^{\prime})\cup R(C^{\prime})\,.$
Near the corner $u_{5}$, the set $g^{0}$ ($\subset D^{\prime}$) intersects $L(C^{\prime})$. Near the corner $u_{3}$, $g^{0}$ intersects $R(C^{\prime})$. Hence $L(C^{\prime})$ and $R(C^{\prime})$ cut $g^{0}$, which is a contradiction because $g^{0}$ is connected by Proposition 1.13. In each of the five other cases (when two $e_{i}^{0}$ lie in $L(C^{\prime})$ or in $R(C^{\prime})$) we obtain a similar contradiction.
Claim 3. $K_{3,3}$-configurations do not exist. We assume that $U$, $V$, and $\{f_{u,v}\colon\;u\in U,v\in V\}$ is a $K_{3,3}$-configuration. We transform it in a PL $K_{3,3}$-configuration, which by Claim 2 is a contradiction. Let $U=\{u_{1},u_{2},u_{3}\}$ and $V=\{u_{4},u_{5},u_{6}\}$. We again write $f_{i,j}$ instead of $f_{u_{i},u_{j}}$. For $i=1,2,\dots,6$, let
$B_{i}=\overline{B}(u_{i},\,r)=\{x\in\mathbb{R}^{2}\colon\;d_{2}(x,\,u_{i})\leq r\}$
be mutually disjoint closed balls with the radius $r>0$ so small that $f_{i,j}[I]\cap B_{l}=\emptyset$ whenever $l\neq i,j$ (Propositions 1.6 and 1.10). Let $D=\mathbb{R}^{2}\setminus\bigcup_{i=1}^{6}B_{i}$. Using the arcs $f_{i,j}$ and Proposition 1.9, we obtain nine subarcs $g_{i,j}\colon I=[a,b]\to\mathbb{R}^{2}$ of $f_{i,j}$ such that $g_{i,j}$ joins a point $\alpha_{i,j}\in\partial B_{i}$ to a point $\beta_{i,j}\in\partial B_{j}$, that $g_{i,j}^{0}\subset D$ and that the nine sets $g_{i,j}[I]$ are mutually disjoint. Using Propositions 1.6, 1.10, and 1.12, we obtain $\delta>0$ and nine partitions of $[a,b]$,
$a=a_{0,\,i,\,j}<a_{1,\,i,\,j}<\dots<a_{k(i,\,j),\,i,\,j}=b\ \ (k(i,\,j)\in\mathbb{N}=\{1,2,\dots\})\,$,
such that
$g_{i,\,j}(a_{l-1,\,i,\,j})\in B(g_{i,\,j}(a_{l,\,i,\,j}),\,\delta)\,\text{ for }\,l=1,\,2,\,\dots,\,k(i,\,j)\,$,
and that
$B(g_{i,\,j}(a_{l,\,i,\,j}),\,\delta)\cap B(g_{i^{\prime},\,j^{\prime}}(a_{l^{\prime},\,i^{\prime},\,j^{\prime}}),\,\delta)=\emptyset$
for every $l=0,1,\dots,k(i,j)$, $l^{\prime}=0,1,\dots,k(i^{\prime},j^{\prime})$ and pairs $\langle i,j\rangle\neq\langle i^{\prime},j^{\prime}\rangle$. Let
$U_{i,\,j}=\bigcup_{l=0}^{k(i,\,j)}B(g_{i,\,j}(a_{l,\,i,\,j}),\,\delta)\,.$}
Each of these nine open sets is connected by Propositions 1.15 and 1.16 (the intersection graph spans a path), and they are mutually disjoint. Using Proposition 1.18, we obtain nine PL arcs $h_{i,j}\colon[a,b]\to U_{i,j}$ such that $h_{i,j}$ joins $\alpha_{i,j}=h_{i,j}(a)=g_{i,j}(a)$ ($\in\partial B_{i}$) to $\beta_{i,j}=h_{i,j}(b)=g_{i,j}(b)$ ($\in\partial B_{j}$) in $U_{i,j}$. We consider the first intersections of the straight segments $u_{i}\,\alpha_{i,j}$ and $u_{j}\,\beta_{i,j}$ with $h_{i,j}[I]$ (Proposition 1.9) and obtain nine PL arcs $p_{i,j}$ such that $p_{i,j}$ joins $u_{i}$ to $u_{j}$ and
$p_{i,\,j}^{0}\cap p_{i^{\prime},\,j^{\prime}}^{0}=\emptyset=p_{i,\,j}^{0}\cap(U\cup V)$
for every pairs $\langle i,j\rangle\neq\langle i^{\prime},j^{\prime}\rangle$. Then $U$, $V$, and $\{p_{i,j}\colon\;1\leq i\leq 3,\,4\leq j\leq 6\}$ is a PL $K_{3,3}$-configuration.
[…]
So let $f\colon I\to\mathbb{R}^{2}$ be a circuit with the property that the complement $D=\mathbb{R}^{2}\setminus f[I]$ is connected. Let $a$ and $b$ be points in $f[I]$ with the maximum and minimum $y$-coordinate, respectively (Propositions 1.6 and 1.19). Clearly, $a_{y}>b_{y}$ (Proposition 1.20). By Proposition 1.21 the points $a$ and $b$ divide $f[I]$ in two arcs $f_{0},f_{1}\colon I\to\mathbb{R}^{2}$ such that $f_{0}^{0}\cap f_{1}^{0}=\emptyset$, $f_{0}[I]\cup f_{1}[I]=f[I]$, and $\mathrm{ep}(f_{0})=\mathrm{ep}(f_{1})=\{a,b\}$. Let $f_{2}\colon I\to\mathbb{R}^{2}$ be a PL arc joining $a$ to $b$ such that $f_{2}^{0}\subset D$. For example, $f_{2}$ starts with a vertical segment going up from $a$, then goes horizontally sufficiently far to the right, goes vertically down to a level below $b_{y}$, continues horizontally left just below $b$, and closes by going vertically up to $b$. Such a PL arc $f_{2}$ exists by Propositions 1.6 and 1.7. Let $\ell$ be any horizontal line with $y$-coordinate in $(b_{y},a_{y})$. By Propositions 1.22 and 1.23, the line $\ell$ contains a straight segment $T$ such that one endpoint $c$ is in $f_{0}^{0}$, the other endpoint $d$ is in $f_{1}^{0}$ and $T^{0}\subset D$. By the connectedness of $D$ and by Propositions 1.8, 1.9 and 1.18, there exists an arc $f_{e,g}$ joining a point $e\in T^{0}$ to a point $g\in f_{2}^{0}$ in $D$ and such that $f_{e,g}^{0}\cap(T\cup f_{2}[I])=\emptyset$. We set $U=\{a,b,e\}$, $V=\{c,d,g\}$ and define the connecting arcs. The arc $f_{e,g}$ is already defined. The arcs $f_{e,c}$ and $f_{e,d}$ are the corresponding subsegments of $T$. The arcs $f_{a,c}$ and $f_{a,d}$ are the corresponding subarcs of $f_{0}$ and $f_{1}$, respectively. The arc $f_{a,g}$ is the corresponding subarc of $f_{2}$. The arcs $f_{b,c}$, $f_{b,d}$, and $f_{b,g}$ are defined similarly to those starting at $a$. It is easy to check that the nine arcs defined have mutually disjoint interiors and that $U$, $V$, and $\{f_{u,v}\colon\;u\in U,v\in V\}$ is a $K_{3,3}$-configuration.
3 Filippov’s proof of the WJT
Step 1. The parity map $N$ is constant. Let $f$ and $g$ be as described in Section 1: $f,g\colon I=[a,b]\to\mathbb{R}^{2}$ are PL maps such that $f[I]\cap g[I]=\emptyset$ and that either $f$ is closed or $g[I]$ lies between two vertical lines going through the points $f(a)$ and $f(b)$. Let $S(f)=\mathbb{R}^{2}\setminus f[I]$ if $f$ is closed, and otherwise let $S(f)$ be the intersection of $\mathbb{R}^{2}\setminus f[I]$ with the open strip between the two lines.
We define the parity map
$N=N(c)=N(c,\,f)\colon S(f)\to\{0,\,1\}\ \ (\subset\mathbb{R})\,.$
Let $c\in S(f)$. As in Claim 1 of Thomassen’s proof, we denote by $\ell(c)$ the half-line going up from $c$ and consider the geometric intersection $P(c)$ of $\ell(c)$ and $f[I]$. For $P(c)=\emptyset$ we set $N(c)=0$. Let $P(c)\neq\emptyset$. Since now $f$ need not be injective, we consider instead of $\ell(c)\cap f[I]$ the preimage of this set. Let $a=a_{0}<a_{1}<\dots<a_{k}=b$ be the partition of the domain $[a,b]$ of $f$. For every $i=1,2,\dots,k$, the segment $s_{i}=f[\,[a_{i-1},a_{i}]\,]$ of $f[I]$ is either disjoint to $\ell(c)$ or $s_{i}\cap\ell(c)=\{p\}$ for a single point $p$ or $s_{i}\subset\ell(c)$. It follows that
$P(c)=P(c,\,f):=f^{-1}[\ell(c)\cap f[I]\,]\ \ (\subset[a,\,b])$
is a finite disjoint union of components $z$ such that every $z$ is either a subinterval $a_{i},a_{j}]$ ($\subset[a,b]$) with $0\leq i<j\leq k$, where for non-closed $f$ the first and last inequality is strict, or a point in $(a,b)$. We call zzsimple component if the values $f(t)$ for $t$ immediately before and after $z$ lie on different sides of $\ell(c)$. Else, if they lie on the same side of $\ell(c)$, we call zzdouble component. We define
[…]
We prove that the parity map $N$ is continuous on $S(f)$. Let $c\in S(f)$. We take a $\delta>0$ small enough such that the ball $B=B(c,\delta)\subset S(f)$ (Proposition 1.8) and that no vertical line that intersects $B$ contains a corner of $f[I]$ distinct from those whose preimages appear in $P(c)$. Let $c^{\prime}\in B$ with $c^{\prime}\neq c$. As we move $c$ to $c^{\prime}$, every simple component of $P(c)$ is preserved and every double component is either fixed or disappears or decays in two simple (point) components. It follows that $N(c)=N(c^{\prime})$. So $N$ is continuous at $c$ and on $S(f)$.
We obtain the main result of Step 1:
[…]
Indeed, since $g[I]\subset S(f)$, the map $N(g)\colon I\to\mathbb{R}$ is continuous and $N(g)[I]\subset\{0,1\}$. If $N(g)[I]=\{0,1\}$, we contradict Proposition 1.13.
Step 2. Two properties of the parity map $N$. The first property is the additivity of $N(c,f)$ in the variable $f$. Let $f_{1},f_{2}\colon I=[a,b]\to\mathbb{R}^{2}$ be two non-closed PL maps such that $f_{1}(b)=f_{2}(a)$. Let $f_{3}=f_{1}+f_{2}$ and let $c\in S(f_{1})\cap S(f_{2})\cap S(f_{3})$. Then
$N(c,\,f_{3})=N(c,\,f_{1})+N(c,\,f_{2})\ \ (\mathrm{mod}\;2)\,.$
Here is an outline of the proof of this equality: if $S_{i}$ is the set of simple components in $P(c,f_{i})$, $i=1,2,3$, then there is a bijection between
$S_{3}\,\text{ and }\,(S_{1}\times\{0\})\cup(S_{2}\times\{1\})\,.$
We leave details of the bijection to the reader as an exercise.
We proceed to the second property of $N$. Let $f\colon I=[a,b]\to\mathbb{R}^{2}$ be a non-closed PL map and let $c\in S(f)$ lie below all intersections of the line $x=c_{x}$ with $f[I]$. Then
N(c,f)=1.N(c,\,f)=1\,.
To prove it, we consider the partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of the domain of $f$ and the components $z_{1}<z_{2}<\dots<z_{l}$ of
$P(c,f)=f^{-1}[\ell(c)\cap f[I]\,]=f^{-1}[(x=c_{x})\cap f[I]\,]\,.$
As we know, every component $z_{i}$ is an interval $a_{r_{i}},a_{s_{i}}]$ ($\subset[a,b]$) with $0<r_{i}<s_{i}<k$ or a point $\{t_{i}\}$ with $t_{i}\in(a,b)$. It is clear that $P(c,f)\neq\emptyset$ and $l\geq 1$ (Proposition 1.22). We call the $l+1$ open subintervals
$g_{0}=(a_{0},\min(z_{1})),\,g_{1}=(\max(z_{1}),\min(z_{2})),\,\dots,\,g_{l}=(\max(z_{l}),\,,a_{k})$
of $(a,b)$ gaps. It follows from the assumption on $c$ that for every gap $g_{i}$ and every $t\in g_{i}$, all values $f(t)$ lie on only one side of $\ell(c)$, the left side $L$ or the right side $R$. We accordingly assign to $P(c,f)$ a word $w_{0}w_{1}\dots w_{l}\in\{L,R\}^{l+1}$. For $i=1,2,\dots,l$, the transition $w_{i-1}w_{i}$ with $w_{i-1}=w_{i}$ corresponds to a double component of $P(c,f)$, and that with $w_{i-1}\neq w_{i}$ to a simple component. Since $w_{0}\neq w_{l}$, the number of the latter transitions is odd.
Step 3. Some two points in the complement of any circuit are separated. Let $f\colon I\to\mathbb{R}^{2}$ be a circuit and $D=\mathbb{R}^{2}\setminus f[I]$. We may assume that $f[I]$ lies to the right of the $y$-axis. Let $l\in f[I]$ and $p\in f[I]$ be, respectively, a leftmost and a rightmost point in $f[I]$ (Propositions 1.6 1.19). These points divide $f$ in two arcs $f_{1},f_{2}\colon I\to\mathbb{R}^{2}$ such that $f_{1}^{0}\cap f_{2}^{0}=\emptyset$, $f_{1}[I]\cup f_{2}[I]=f[I]$ and $\mathrm{ep}(f_{1})=\mathrm{ep}(f_{2})=\{l,p\}$ (Proposition 1.21). Clearly, $l_{x}<p_{x}$ (Proposition 1.20). Let $\ell$ be any vertical line with $x$-coordinate in $(l_{x},p_{x})$. Let $a\in f[I]\cap\ell$ be the highest point of this intersection (Proposition 1.19). We may assume that $a\in f_{1}^{0}$. Similarly, let $b\in f_{1}^{0}\cap\ell$ be the lowest point of this intersection. We take a point $c\in\ell$ below $b$ but such that
$0<d_{2}(c,\,b)<d_{2}(b,f_{2}[I])$
— then $c\in D$. The point $c$ exists by Propositions 1.6 and 1.10, by which $d_{2}(b,f_{2}[I])>0$. We show that for every PL map $g\colon I\to\mathbb{R}^{2}$ that joins $c\in D$ to the origin $\overline{0}=\langle 0,0\rangle\in D$ ($f[I]$ lies to the right of the $y$-axis), $g[I]\cap f[I]\neq\emptyset$. In view of Propositions 1.8 and 1.18 it means that the complement $D=\mathbb{R}^{2}\setminus f[I]$ is disconnected.
Suppose for the contrary that $g\colon I\to\mathbb{R}^{2}$ is a PL map joining $c$ to $\overline{0}$ in $D$, so that $g[I]\cap f[I]=\emptyset$. Let $\kappa$ ($\subset\mathbb{R}^{2}$) be the half-line going down from $c$ and let
$\lambda=c\,b\cup f_{3}[I]\cup\ell(a)\ \ (\subset\mathbb{R}^{2})\,$,
where $cb$ is the straight segment joining these two points, $f_{3}$ is a subarc of $f_{1}$ joining $b$ to $a$ (if $b=a$, we omit $f_{3}[I]$) and, as before, $\ell(a)$ is the half-line going up from $a$. Let $h>0$ be the minimum of the three positive distances
$d_{2}(g[I],\,f[I])>0,\ d_{2}(\kappa,\,f_{1}[I])>0\,\text{ and }\,d_{2}(\lambda,\,f_{2}[I])>0\,.$
The first distance is positive by the assumption that $g[I]\cap f[I]=\emptyset$ and by Propositions 1.6 and 1.10. The latter two distances are positive by Propositions 1.4, 1.5, 1.6, 1.7, and 1.11.
We approximate the given circuit $f\colon I=[\alpha,\beta]\to\mathbb{R}^{2}$ sufficiently closely by a closed PL map. By Proposition 1.12, there exists a partition $\alpha=a_{0}<a_{1}<\dots<a_{k}=\beta$ of $\alpha,\beta]$ such that $\{a,b,l,p\}\subset\{f(a_{i})\colon\;i=0,1,\dots,k\}$ and
$d_{2}(f(a_{i-1}),\,f(a_{i}))<h\,\text{ for every }\,i=1,\,2,\,\dots,\,k\,.$
We join for $i=1,2,\dots,k$ the consecutive points $f(a_{i-1})$ and $f(a_{i})$ ($f(a_{0})=f(a_{k})$) by straight segments and obtain a closed PL map $f_{4}\colon I=[\alpha,\beta]\to\mathbb{R}^{2}$ such that all corners of $f_{4}[I]$ lie on $f[I]$ and include the four points $a$, $b$, $l$, and $p$. Also, for every $z\in f_{4}[I]$ we have $d_{2}(z,f[I])<\frac{1}{2}h$. Let $f_{5}$, $f_{6}$, and $f_{7}$ be the non-closed PL sub-maps of $f_{4}$ spanned by the corners of $f_{4}[I]$ lying in $f_{1}[I]$, $f_{2}[I]$, and $f_{3}[I]$, respectively. The above definition of $h$ and the $\frac{1}{2}h$-approximation of $f_{4}[I]$ by $f[I]$ imply that
$f_{4}[I]\cap g[I]=\emptyset,\ f_{5}[I]\cap\kappa=\emptyset,\,\text{ and }\,f_{6}[I]\cap(cb\cup f_{7}[I]\cup\ell(a))=\emptyset\,.$
The parity function $N$ produces the contradiction
$0=N(\overline{0},\,f_{4})=N(c,\,f_{4})=N(c,\,f_{5})+N(c,\,f_{6})=1+0=1\,.$
The first equality follows from the fact that since $f[I]$ lies to the right of the $y$-axis, so does $f_{4}[I]$. The second equality follows from Step 1 because $f_{4}[I]\cap g[I]=\emptyset$. The third equality follows from the first property of $N$ in Step 2 because (we can define $f_{5}$ and $f_{6}$ in such a way that) $f_{4}=f_{5}+f_{6}$. We have $N(c,f_{5})=1$ by the second property of $N$ in Step 2 (since $f_{5}[I]\cap\kappa=\emptyset$, the point $c$ lies below all intersections of the line $x=c_{x}$ with $f_{5}[I]$). Finally, $N(c,f_{6})=N(a,f_{6})$ by Step 1 because $(cb\cup f_{7}[I])\cap f_{6}[I]=\emptyset$, and $N(a,f_{6})=0$ by the definition of $N$ because $\ell(a)\cap f_{6}[I]=\emptyset$.
4 Concluding remarks
Jordan’s proof of his theorem in [14] was often questioned in the past, and it was claimed that the first correct proof of the JT was given by Veblen [18]. Hales [8] reversed these assessments and showed that Jordan’s proof is basically correct when some gaps are filled, but that Veblen’s proof has serious problems. Filippov’s article [5] is preceded by the article [19] by A. I. Vol’pert, which contains a similar proof of the JT. The footnote on the first page of [5] reads “After this work was completed, the author learned that a proof of Jordan’s theorem, very close to the one in this article, was somewhat earlier found by A. I. Vol’pert (Lviv)”. Authors of these two proofs, Aleksei F. Filippov (1923–2006) and Aizik I. Volpert (Vol’pert) (1923–2006), were excellent mathematicians — see [2, 1].
In recent years, the ways in which we (mathematicians and philosophers) think about the links between informal and formal proofs, and about the role of formalized mathematics, have changed considerably and dramatically compared to the times before, say, 2000: see, for example, [3, 4, 6, 7]. What should be the results of the work of a research mathematician? A possible answer is: rigorous proofs of mathematical theorems. An informal proof is rigorous if it can be routinely translated into a formal proof ([11]), formal in the sense of the previously mentioned developments; see also [9, 10]. In this article, we wanted to lift the beautiful proofs of Thomassen and Filippov to the modern level of rigor, and we hopefully succeeded. The preprint [3] reports on the process of autoformalization of Munkres’ textbook [16] on topology. It is an impressive achievement; it seems that very soon AI will do everything for us. But does this autoformalization help a student interested in topology learn and understand ([12, 13]) the proofs of theorems in [16]? Not much. In contrast, we hope that our article will be helpful to anyone wishing to understand the two proofs of the WJT.
I lectured on Thomassen’s proof of the WJT in the course Mathematical Analysis 3 in the years 2023–2026. For the next year, I plan to switch to the proofs of the JT in [5, 19].